Find the current price of a one-year, R110-strike American put option on a non-
dividend-paying stock whose current price is S(0) = 100. Assume that the continuously compounded interest rate equals r = 0.06. Use a two-period Binomial tree with
u = 1.23, and d = 0.86 to calculate the price VP(0) of the put option.
1
Expert's answer
2021-07-27T02:15:01-0400

"P=\frac{1+r-d}{U-d}\\\frac{1+0.06-0.86}{1.23-0.86}\\\frac{0.2}{0.37}=54\%"

OP of put today

"\frac{11.04}{e^{0.06×1}}=10.397
\\=e^{0.06}=1.0618"